OpenStudy (idealist10):

Find the orthogonal trajectories of y=ce^(2x).

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

I differentiated y=ce^(2x) with respect to x and got dy/dx=2ce^(2x). So the negative reciprocal of dy/dx=-1/(2ce^(2x)) and I need to solve for y from here.

OpenStudy (idealist10):

Oh yeah, isn't it dy/dx=-1/2y?

OpenStudy (anonymous):

Okay, so you have the separable equation, \[\frac{dy}{dx}=-\frac{1}{2c}e^{-2x}\] Given that \(y=ce^{2x}\), you have \(c=ye^{-2x}\), so yes, you get \[\frac{dy}{dx}=-\frac{1}{2y}\]

OpenStudy (idealist10):

I was thoughtless. See, I was stucked because I thought there's no y to replace it and substitute. Now I got it. Thanks for your time!